OpenStudy (anonymous):

Use the distributive property to rewrite the expression and solve. 6 * 94 = 6 * ? + 6 4 = ?

OpenStudy (anonymous):

@perl

OpenStudy (anonymous):

@Princer_Jones

OpenStudy (anonymous):

Use the distributive property to rewrite the expression and solve. 6 * 94 = 6 * ? + 6 * 4 = ?

OpenStudy (anonymous):

6(94)=6(90)+6(4)=540+24=564

OpenStudy (anonymous):

@Daveqwerty

OpenStudy (anonymous):

since 94=90+4

OpenStudy (anonymous):

@Daveqwerty solution clear to you?